Attempts: Average: 12 Attention: Due to a bug in Google Chrome, this page may not function correcty, Click here to learn more 12. Equivalent annual annuities Another method to deal with the unequal life problem of projects is the equivalent annual annuity (EAA) method. In this method the annual cash flows under the alternative investments are converted into a constant cash flow stream whose NPV is equivalent to the NPV of the comparative projects initial stream. Consider the case of Cold Duck Manufacturing Company Cold Duck Manufacturing Company is considering a four-year project that has a weighted average cost of capital of 11% and a net present value (NPV) of $75,682. Cold Duck Manufacturing Company can replicate this project indefinitely. The equivalent annual annuity (EAA) for this project is do a good job of taking inflation into The EAA approach to evaluating projects with unequal lives account. Flash Playor WIN 32.0,0,114 03 3.34.1 2004-2016 Ap All rights 2013 C
Answer #1

Equivalent annual annuity = NPV/ (P/A , 11%,4)

= 75682/ (1-(1/(1+r)^n)/r)

= 75682/ (1-1/1.11^4)/0.11

=$24394.3

b: DOES NOT

The EAA approach does not consider the impact of inflation.